OpenStudy (hulagirlmb):

A chemist wishes to make 3 liters of a 7% acid solution by mixing a 9% acid solution and a 4% acid solution. Let x represent the number of liters of the 9% acid solution. Write an equation to solve for x.

OpenStudy (hulagirlmb):

I only have 15 minutes before it's due! Please help

OpenStudy (anonymous):

ok lets work quick

OpenStudy (hulagirlmb):

I'll post the answer choices in a drawing one second

OpenStudy (anonymous):

\(x\) is the number of liters of 9%k solution, making the number of liters of 4% solutions \(3-x\) solve \[.09x+.04(3-x)=.07\times 3\]

OpenStudy (hulagirlmb):

Thank you

OpenStudy (anonymous):

or better still, solve \[9x+4(3-x)=7\times 3\]

OpenStudy (hulagirlmb):

Ok one sec

OpenStudy (anonymous):

\[9x+12-4x=21\\ 5x+12=21\\ 5x=9\\ x=\frac{9}{5}=1.8\]

OpenStudy (hulagirlmb):

I got x=1.8

OpenStudy (hulagirlmb):

Ok good

OpenStudy (anonymous):

yay

OpenStudy (hulagirlmb):

Thank you so much, I have to get a good grade on this XD

OpenStudy (anonymous):

good luck that all or you got more?

OpenStudy (hulagirlmb):

I already submitted, somebody else helped me on 2 of the problems and we didn't get them right. But thank you so much for helping me on those 2 because I really didn't understand that type of problem

OpenStudy (anonymous):

i don't think i skipped any steps, you can look back and see if they are clear and ask if they are not this 'mixture" problems are usually confusing at first

OpenStudy (hulagirlmb):

Ok thank you

OpenStudy (perl):

nice work